# On a conjecture concerning the property of chromatic polynomials with negative variable

### 08 Mar 2026 v1

@@ Abstract
Let $G$ be a graph of order $n$ and $P(G,x)$ be the chromatic polynomial of $G$. Dong, Ge, Gong, Ning, Ouyang, and Tay (J. Graph Theory 96 (2021) 343) conjectured that $\frac{d^{k}}{dx^{k}}\bigl(\ln[(-1)^{n}P(G,x)]\bigr)<0$ holds for all $k\geq 2$ and $x\in(-\infty,0)$. We prove this conjecture for all $k\geq 2$ and $x\leq-10\Delta k$, in which $\Delta$ is the maximum degree of $G$.
@@ 1 Introduction
There exists a constant $K$ such that for every graph $G$, $P(G,x)\neq 0$ for all $x\in\mathbb{C}$ with $|x|\geq K\Delta$.
Sokal [10] showed the constant $K\approx 7.97$, Fernández and Procacci [6] improved to $K\approx 6.91$, Jenssen, Patel, Regts [8] recently showed $K\leq 5.94$.
@@ 2 Main results
Because $K\leq 5.94$, $|c_{i}|\leq\frac{n}{i}(K\Delta)^{i}<\frac{n}{i}(6\Delta)^{i}$. And
@@
For $x\leq q$, $q<-6\Delta$ and $k\geq 2$, when $i$ is odd, $\frac{(-1)^{k}}{x^{i+k}}<0$,
$(c_{i}x^{-i})^{(k)}\leq\dfrac{-n(6\Delta)^{i}(-1)^{k}(i+1)\cdots(i+k-1)}{x^{i+k}}=n(6\Delta)^{i}(\frac{1}{x^{i+1}})^{(k-1)};$
@@
$(c_{i}x^{-i})^{(k)}\leq\dfrac{n(6\Delta)^{i}(-1)^{k}(i+1)\cdots(i+k-1)}{x^{i+k}}=-n(6\Delta)^{i}(\frac{1}{x^{i+1}})^{(k-1)}.$
@@
$\leq$ $\frac{n}{6\Delta}(\sum_{\begin{subarray}{c}i=3\\ i\text{ is odd}\end{subarray}}^{\infty}(\frac{6\Delta}{x})^{i+1})^{(k-1)}-\frac{n}{6\Delta}(\sum_{\begin{subarray}{c}i=4\\ i\text{ is even}\end{subarray}}^{\infty}(\frac{6\Delta}{x})^{i+1})^{(k-1)}$
$=$ $\frac{n}{6\Delta}(\frac{(6\Delta)^{4}}{x^{3}(x+6\Delta)})^{(k-1)}=n(\frac{1}{x}-\frac{6\Delta}{x^{2}}+\frac{(6\Delta)^{2}}{x^{3}}-\frac{1}{x+6\Delta})^{(k-1)}$
$=$ $n(-1)^{k-1}\Big(\frac{(k-1)!}{x^{k}}-6\Delta\frac{k!}{x^{k+1}}+(6\Delta)^{2}\frac{(k+1)!}{2x^{k+2}}-\frac{(k-1)!}{(x+6\Delta)^{k}}\Big)$
@@
$+n(-1)^{k-1}\Big(\frac{2(k-1)!}{x^{k}}-6\Delta\frac{k!}{x^{k+1}}+(6\Delta)^{2}\frac{(k+1)!}{2x^{k+2}}-\frac{(k-1)!}{(x+6\Delta)^{k}}\Big)$
$=$ $n(-1)^{k-1}(k-1)!\frac{(x+6\Delta)^{k}\big(8x^{2}-22x\Delta k+k(k+1)+2(6\Delta)^{2}k(k+1)\big)-4x^{k+2}}{4x^{k+2}(x+6\Delta)^{k}}$
To prove $(\ln[(-1)^{n}P(G,x)])^{(k)}<0$ for $x<-10\Delta k$, it is enough to verify the following inequality for $x<-10\Delta k$,
$(-1)^{k-1}(x+6\Delta)^{k}\big(8x^{2}-22x\Delta k+k(k+1)+2(6\Delta)^{2}k(k+1)\big)<(-1)^{k-1}4x^{k+2}.$ (7)
We denote $f_{1}(x)=8x^{2}-22x\Delta k+k(k+1)+2(6\Delta)^{2}k(k+1)$. Because $f_{1}(x)>0$ and $(-1)^{k-1}x^{k}<0$, inequality (7) is equivalent to
$\frac{(-1)^{k-1}(x+6\Delta)^{k}}{(-1)^{k-1}x^{k}}>\frac{4x^{2}}{f_{1}(x)},$
@@
$(1+\frac{6\Delta}{x})^{k}>\frac{4x^{2}}{f_{1}(x)}.$
By using Bernoulli’s Inequality, we have $(1+\frac{6\Delta}{x})^{k}\geq 1+\frac{6\Delta k}{x}.$ We let
$1+\frac{6\Delta k}{x}>\frac{4x^{2}}{f_{1}(x)}$
@@
$xf_{1}(x)+6k\Delta f_{1}(x)-4x^{3}<0.$
Let $f(x)=xf_{1}(x)+6k\Delta f_{1}(x)-4x^{3}.$ We can check that when $x\leq-6\Delta k$, $f^{\prime}(x)>0$, $f(x)$ is increasing strictly. By computing, we have $f(-10\Delta k)<0$. Hence when $x\leq-10\Delta k$, $f(x)<0$, $(\ln[(-1)^{n}P(G,x)])^{(k)}<0$ follows. The proof is completed. ∎
@@ Acknowledgements
The author is supported by National Natural Science Foundation of China (No. 12371350).

### 16 Mar 2026 v2

@@ Abstract
Let $G$ be a graph of order $n$ and $P(G,x)$ be the chromatic polynomial of $G$. Dong, Ge, Gong, Ning, Ouyang, and Tay (J. Graph Theory 96 (2021) 343) conjectured that $\frac{d^{k}}{dx^{k}}\bigl(\ln[(-1)^{n}P(G,x)]\bigr)<0$ holds for all $k\geq 2$ and $x\in(-\infty,0)$. We prove this conjecture for all $k\geq 2$ and $x\leq-6.66\Delta k$, in which $\Delta$ is the maximum degree of $G$.
@@ 1 Introduction
There exists a constant $K$ such that for every graph $G$, $P(G,x)\neq 0$ for all $x\in\mathbb{C}$ with $|x|>K\Delta$.
Sokal [11] showed the constant $K\leq 7.97$, Fernández and Procacci [7] improved to $K\leq 6.91$, Jenssen, Patel, Regts [9] further improved to $K\leq 5.94$. Very recently, Bencs and Regts [3] established a better bound $K\leq 4.25$.
@@ 2 Main results
Because $K\leq 4.25$, $|c_{i}|\leq\frac{n}{i}(K\Delta)^{i}\leq\frac{n}{i}(4.25\Delta)^{i}$. And
@@
For $x\leq q$, $q<-4.25\Delta$ and $k\geq 2$, when $i$ is odd, $\frac{(-1)^{k}}{x^{i+k}}<0$,
$(c_{i}x^{-i})^{(k)}\leq\dfrac{-n(4.25\Delta)^{i}(-1)^{k}(i+1)\cdots(i+k-1)}{x^{i+k}}=n(4.25\Delta)^{i}(\frac{1}{x^{i+1}})^{(k-1)};$
@@
$(c_{i}x^{-i})^{(k)}\leq\dfrac{n(4.25\Delta)^{i}(-1)^{k}(i+1)\cdots(i+k-1)}{x^{i+k}}=-n(4.25\Delta)^{i}(\frac{1}{x^{i+1}})^{(k-1)}.$
@@
$\leq$ $\frac{n}{4.25\Delta}(\sum_{\begin{subarray}{c}i=3\\ i\text{ is odd}\end{subarray}}^{\infty}(\frac{4.25\Delta}{x})^{i+1})^{(k-1)}-\frac{n}{4.25\Delta}(\sum_{\begin{subarray}{c}i=4\\ i\text{ is even}\end{subarray}}^{\infty}(\frac{4.25\Delta}{x})^{i+1})^{(k-1)}$
$=$ $\frac{n}{4.25\Delta}(\frac{(4.25\Delta)^{4}}{x^{3}(x+4.25\Delta)})^{(k-1)}=n(\frac{1}{x}-\frac{4.25\Delta}{x^{2}}+\frac{(4.25\Delta)^{2}}{x^{3}}-\frac{1}{x+4.25\Delta})^{(k-1)}$
$=$ $n(-1)^{k-1}\Big(\frac{(k-1)!}{x^{k}}-4.25\Delta\frac{k!}{x^{k+1}}+(4.25\Delta)^{2}\frac{(k+1)!}{2x^{k+2}}-\frac{(k-1)!}{(x+4.25\Delta)^{k}}\Big)$
@@
$+n(-1)^{k-1}\Big(\frac{2(k-1)!}{x^{k}}-4.25\Delta\frac{k!}{x^{k+1}}+(4.25\Delta)^{2}\frac{(k+1)!}{2x^{k+2}}-\frac{(k-1)!}{(x+4.25\Delta)^{k}}\Big)$
$=$ $n(-1)^{k-1}(k-1)!\frac{(x+4.25\Delta)^{k}\big(8x^{2}-15x\Delta k+k(k+1)+2(4.25\Delta)^{2}k(k+1)\big)-4x^{k+2}}{4x^{k+2}(x+4.25\Delta)^{k}}$
To prove $(\ln[(-1)^{n}P(G,x)])^{(k)}<0$ for $x\leq-6.66\Delta k$, it is enough to verify the following inequality for $x\leq-6.66\Delta k$,
$(-1)^{k-1}(x+4.25\Delta)^{k}\big(8x^{2}-15x\Delta k+k(k+1)+2(4.25\Delta)^{2}k(k+1)\big)<(-1)^{k-1}4x^{k+2}.$ (7)
We denote $f_{1}(x)=8x^{2}-15x\Delta k+k(k+1)+2(4.25\Delta)^{2}k(k+1)$. Because $f_{1}(x)>0$ and $(-1)^{k-1}x^{k}<0$, inequality (7) is equivalent to
$\frac{(-1)^{k-1}(x+4.25\Delta)^{k}}{(-1)^{k-1}x^{k}}>\frac{4x^{2}}{f_{1}(x)},$
@@
$(1+\frac{4.25\Delta}{x})^{k}>\frac{4x^{2}}{f_{1}(x)}.$
By using Bernoulli’s Inequality, we have $(1+\frac{4.25\Delta}{x})^{k}\geq 1+\frac{4.25\Delta k}{x},$ when $x<-4.25\Delta k$. We let
$1+\frac{4.25\Delta k}{x}>\frac{4x^{2}}{f_{1}(x)}$
@@
$xf_{1}(x)+4.25k\Delta f_{1}(x)-4x^{3}<0.$
Let $f(x)=xf_{1}(x)+4.25k\Delta f_{1}(x)-4x^{3}.$ We can check that when $x<-4.25\Delta k$, $f^{\prime}(x)>0$, $f(x)$ is increasing strictly. By computing, we have $f(-6.66\Delta k)<0$. Hence when $x\leq-6.66\Delta k$, $f(x)<0$, $(\ln[(-1)^{n}P(G,x)])^{(k)}<0$ follows. The proof is complete. ∎
We note that to obtain inequalities (2) and (3) required for the proof of Theorem 2.5, we need to estimate the value of $c_{i}$. In this note, we estimate $c_{i}$ by using $|c_{i}|\leq\frac{n}{i}(K\Delta)^{i}$. Thus, a new bound on $K$ gives a new bound on $c_{i}$, and a smaller $K$ yields a better bound.
@@ Acknowledgements
The author thanks Guus Regts for informing her of reference [3], which provides the best bound on $K$ to date and is very helpful for the present note. The author is supported by National Natural Science Foundation of China (No. 12371350).